OpenStudy (anonymous):

Solve. –3(5 – 15n) = 15(3n – 2)

OpenStudy (anonymous):

\(\Large –3(5 – 15n) = 15(3n – 2)\) \(\Large (-3)\times 5 + (-3)\times (-15n)=15\times 3n + 15\times (-2)\) \(\Large -15+45n=45n-30\) \(\Large -15=-30\) \(\Large \color{red}{\Rightarrow no\ solution}\)

OpenStudy (anonymous):

\(\Large \color{blue}{Method\ 2}\) \(\Large –3(5 – 15n) = 15(3n – 2)\) \(\Large -3\times 5(1-3n)=15(3n-2)\) \(\Large -15(1-3n)=15(3n-2)\) \(\Large 3n-1=3n-2\) \(\Large 1=2\) So, the Equation Has No Solution